1) Consider the population distrībution of X as shown below: f(x) 0.2 0.2 0.6 7 Total Samples of size n-2 will be drawn from this population, by sampling with replacement.
Answer #1

In this Case

Number of possible samples are = 32 = 9

a) Now we make a List of samples with mean and variance

Mean = (\sum X) / n

Variance = \sqrt{\frac{\sum (X - Xbar)^2}{n - 1 }}

List of samples Mean Variance
(1,1) 1 0
(1,7) 4 18
(1,9) 5 32
(7,1) 4 18
(7,7) 7 0
(7,9) 8 2
(9,1) 5 32
(9,7) 8 2
(9,9) 9 0

b) Now construct probability distribution of mean

Mean Probability
1 1/9
4 2/9
5 2/9
7 1/9
8 2/9
9 1/9

c) And probablity distribution of variance

Variance Probability
0 3/9
2 2/9
18 2/9
32 2/9

2) E(Xbar) = ∑x* p(x) = 1 * 1/9 + 4*2/9 + 5*2/9 + 7*1/9 + 8*2/9 + 9*1/9 = 5.667

population mean = \mu = (1 + 7 + 9) / 3 = 5.667

Hence E(xbar) = \mu

As well as

Population standard deviation(\sigma) = \sqrt\frac{\sum(X - \mu)^2}{N} = 3.399

Standard deviation of mean, sd(xbar) = \sqrt{\sum x^2 * p(x) ~-~ \mu^2} = 2.4

  \frac{\sigma}{\sqrt{n}} = \frac{3.399}{\sqrt{2}} = 2.4

Hence SD(Xbar) = \frac{\sigma}{\sqrt{n}}

Please let me know further clarifications regarding this work